One number is equal to the square of another. Find the numbers if both are positive and their sum is 1482.
Answer by TimothyLamb(4379)

y = xx
y + x = 1482
xx + x = 1482
xx + x - 1482 = 0
---
the above quadratic equation is in standard form, with a=1, b=1 and c=-1482
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 1 -1482
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 38
x = -39
---
the negative root doesn't fit the problem statement, so use the positive root:
---
x = 38
y = xx
y = 38*38
y = 1444
---
answer:
x = 38
y = 1444
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php